1. (4 pts) Determine the eigenvalues of 3 -6 А -5 2
Answer #1

Given, A=13 - 6 2 Sol. (A-d I) = 0 3-d. .6 2-d (3-4) (2-) -(-6)(-5) = 0 6-30-2n+d²_3050 2 ? - Sd-24O 2 d -8d + 3d 24-0 - WINin case of doubt please comment